# arXiv:2602.24042v1[cs.DS] 27 Feb 2026

### v1

@@ E.1 Stronger Lower Bound on the Adaptivity Gap
(Terminal item) We say item $i$ is a terminal item if w.p. at least $1-\frac{1}{a}$ it has a size of $2$, for a very large $a\gg n$, and with probability $\frac{1}{a}$ it has a fixed size $s_{i}\in[0,1]$. An item that is not terminal is referred to as a non-terminal item.
@@
We defer the proof of 12 to Section D.3. The idea in the proof of the first property is to fix an instance $I\in\mathcal{H}_{2}$ and replace any item $i\in[n]$ with an “almost” terminal item such that the effective value remains the same, and the size is the original size w.p. $\frac{1}{a}$ or overflows the knapsack (size $2$) otherwise. The second transformation is to replace each such new item with $|C|$ terminal items where $C$ is the support of the size distribution of item 0. In the new resulting instance, all of the items except of item 0 are terminals, and the adaptivity gap may only increase (up to $\frac{1}{a}\ll 1$).
@@ E.2 Adaptivity Gap for $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ Distributions
In this section we show that for $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ item size distributions the adaptivity gap is at most $2$ (rather than the general known gap upper bound of $4$).
@@
The adaptivity gap for the $\operatorname{Non-Risky-SK}$ problem of any instance $I$ consisting of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items is at most $2$: $G(I)\leq 2$.
@@
Starting from an optimal adaptive tree $T$ on $I$, observe that whenever an item realizes to its “large” size, any further adaptive branching is useless (any nonzero size thereafter forces overflow). Thus each such branch of $T$ is a single‐chain of small‐size realizations followed by an eventual stop. Replace each such chain by a single terminal item whose value equals the expected value of that chain. Call the new tree $T^{\prime}$ and the augmented instance $I^{\prime}$, see Fig. 9 for illustration.
Figure 9: An illustration of the reduction from II (with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items only) to $I^{\prime}$ (where the optimal adaptive decision tree is to take some $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items and then a terminal item). On the left there is an illustration of $T$, an optimal adaptive decision tree for instance $I$. On the right there is an illustration of an optimal adaptive decision tree $T^{\prime}$ for instance $I^{\prime}$.
The adaptive reward of $T^{\prime}$ on $I^{\prime}$ equals that of $T$ on $I$, so $\operatorname{ADAPT}(I)\leq\operatorname{ADAPT}(I^{\prime})$. Conversely, any non‐adaptive plan on $I^{\prime}$ that picks a terminal can be simulated in $I$ by following the original chain, so $\operatorname{ALG}(I)\geq\operatorname{ALG}(I^{\prime})$. Hence
@@
We are now ready to show the $2$ upper bound.
@@
Consider a non-adaptive strategy $\operatorname{\sigma_{ALG^{\prime}}}$ that picks the best of these three options. This may not be $\operatorname{\sigma_{ALG}}$ (which might achieve a higher expected value through a different order), but as we will show, bounding the gap between $\operatorname{\sigma_{AD}}$ and $\operatorname{\sigma_{ALG^{\prime}}}$ suffices for proving the upper bound of $2$. Let $\operatorname{ALG}^{\prime}$ denote the expected value of $\operatorname{\sigma_{ALG^{\prime}}}$. Then:

### v2

@@ E.1 Stronger Lower Bound on the Adaptivity Gap
(Terminal item) We say item $i$ is a terminal item if w.p. at least $1-\frac{1}{a}$ it has a size of 22, for a very large $a\gg n$, and with probability $\frac{1}{a}$ it has a fixed size $s_{i}\in[0,1]$. An item that is not terminal is referred to as a non-terminal item.
@@
We defer the proof of 12 to Section D.3. The idea in the proof of the first property is to fix an instance $I\in\mathcal{H}_{2}$ and replace any item $i\in[n]$ with an “almost” terminal item such that the effective value remains the same, and the size is the original size w.p. $\frac{1}{a}$ or overflows the knapsack (size 22) otherwise. The second transformation is to replace each such new item with $|C|$ terminal items where $C$ is the support of the size distribution of item 0. In the new resulting instance, all of the items except of item 0 are terminals, and the adaptivity gap may only increase (up to $\frac{1}{a}\ll 1$).
@@ E.2 Adaptivity Gap for $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ Distributions
In this section we show that for $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ item size distributions the adaptivity gap is at most 22 (rather than the general known gap upper bound of 44).
@@
The adaptivity gap for the $\operatorname{Non-Risky-SK}$ problem of any instance II consisting of $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items is at most 22: $G(I)\leq 2$.
@@
Starting from an optimal adaptive tree $T$ on II, observe that whenever an item realizes to its “large” size, any further adaptive branching is useless (any nonzero size thereafter forces overflow). Thus each such branch of $T$ is a single‐chain of small‐size realizations followed by an eventual stop. Replace each such chain by a single terminal item whose value equals the expected value of that chain. Call the new tree $T^{\prime}$ and the augmented instance $I^{\prime}$, see Fig. 9 for illustration.
Figure 9: An illustration of the reduction from II (with $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items only) to $I^{\prime}$ (where the optimal adaptive decision tree is to take some $\operatorname{\varepsilon-\text{Noisy Bernoulli}}$ items and then a terminal item). On the left there is an illustration of $T$, an optimal adaptive decision tree for instance II. On the right there is an illustration of an optimal adaptive decision tree $T^{\prime}$ for instance $I^{\prime}$.
The adaptive reward of $T^{\prime}$ on $I^{\prime}$ equals that of $T$ on II, so $\operatorname{ADAPT}(I)\leq\operatorname{ADAPT}(I^{\prime})$. Conversely, any non‐adaptive plan on $I^{\prime}$ that picks a terminal can be simulated in II by following the original chain, so $\operatorname{ALG}(I)\geq\operatorname{ALG}(I^{\prime})$. Hence
@@
We are now ready to show the 22 upper bound.
@@
Consider a non-adaptive strategy $\operatorname{\sigma_{ALG^{\prime}}}$ that picks the best of these three options. This may not be $\operatorname{\sigma_{ALG}}$ (which might achieve a higher expected value through a different order), but as we will show, bounding the gap between $\operatorname{\sigma_{AD}}$ and $\operatorname{\sigma_{ALG^{\prime}}}$ suffices for proving the upper bound of 22. Let $\operatorname{ALG}^{\prime}$ denote the expected value of $\operatorname{\sigma_{ALG^{\prime}}}$. Then: